# Unleashing the Potential of Large Language Models as Prompt Optimizers: An Analogical Analysis with Gradient-based Model Optimizers

### v2

@@ A.2 The Impact of Initial Prompts
In our main experiment, we take “Let’s think step by step.” as the initial prompt. In this part, we aim to explore the impact of initial task prompts. Specifically, we consider prompts from four categories (instructive, misleading, irrelevant, and empty) following Kojima et al. (2022). Table 8 presents the results on the lite BBH benchmark. In general, GPO can consistently boost performance across various initial prompts. This observation underscores the versatility of GPO as a prompt optimizer. Furthermore, the efficacy of optimization is more pronounced with relevant initial prompts (i.e., instructive and misleading). It means that prompt initialization is still important, especially in conveying task-specific information.
A.3 The Generalizability of Optimized Prompts
Method Prompt Accuracy
MultiArith AQUA
Zero-shot CoT Let’s think step by step. 58.00 19.69
OPRO
Take a deep breath and work on this problem step-by-step.
59.33 20.47
GPO
Thoroughly strategize and meticulously plan every single
aspect of your approach, leaving no room for oversight,
and give utmost attention to even the smallest details,
with the goal of achieving exceptional results and excelling.
62.33 24.80
To assess the generalizability of optimized prompts, following Zhou et al. (2023), we evaluate the prompts optimized on GSM8k Cobbe et al. (2021) by OPRO and GPO on two additional math reasoning benchmarks MultiArith Roy and Roth (2016) and AQUA Ling et al. (2017). Table 9 demonstrates that the optimized prompts by GPO can be generalized well on these two benchmarks, which perform better than those by OPRO.
A.4 Hyper-Parameter Analysis
@@
A.5 The Execution Time of LLM-based Prompt Optimizers
Figure 8: Execution time of LLM-based prompt optimizers
To further assess the computation costs of LLM-based prompt optimizers, we conduct five repetition experiments and record their average execution time. Figure 8 shows the results on the lite BBH benchmark. We can observe that the execution time is almost consistent with the trend of token consumption (in Section 7). The average execution time of GPO is much lower than SGDM, APO, and PE2, and comparable to OPRO.
Appendix B The Interpretability of Optimized Prompts
@@
Appendix C Additional Details for the Setup of Analogical Analysis
Tasks and Metrics. Following Yang et al. (2023), we utilize Big-Bench Hard (BBH) Suzgun et al. (2023) for evaluation. BBH is a suite of 23 challenging BIG-Bench tasks Srivastava et al. (2022) that covers various kinds of reasoning capabilities. Due to the constraints of computational resources, we select a dataset from each type of task in BBH to create a lite benchmark for our analysis: i) Navigate (Binary choice); ii) Movie Recommendation (Multiple choice); iii) Object Counting (Numeric response); iv) Word Sorting (Free response). For each dataset, we split it into train/valid/test sets with a ratio of 2:2:6. Following Suzgun et al. (2023), we adopt the exact match as the metric for performance evaluation.
Implementation Details. Following Crispino et al. (2023), we select Llama-2-7b-chat Touvron et al. (2023) as the task model and set its temperature to 0 to make the output as deterministic as possible. For the prompt optimizer, we employ gpt-3.5-turbo, which is the underlying model of ChatGPT. We set its temperature to 1.0 to encourage the generation to be more diverse. To help the prompt optimizer understand the downstream task, following Yang et al. (2023), we randomly sample 3 examples from the dataset and fill them into the meta-prompt of the prompt optimizer. In the process of optimization, we take the instruction “Let’s think step by step.” as the initial prompt and insert it at the end of the question to obtain better performance following Suzgun et al. (2023). At each step, we first prompt the optimizer to generate 8 candidate task prompts, and then select the best-performing one as the task prompt for the next iteration following Yang et al. (2023); Ye et al. (2023). The optimization process lasts for at most 3 epochs with a batch size of 8. We repeat all the experiments three times and report the average of the results. The meta-prompts we used are detailed in Appendix E.1.
Appendix D Additional Details for the Setup of Experiment
D.1 The Statistics of Datasets
As mentioned in Section 4.1, we sample a subset of the dataset for efficient evaluation. For BBH and MMLU, we split the entire dataset into training, validation, and test sets with a ratio of 2:2:6. For GSM8k and WebNLG, we randomly sample 100 examples as the training set, 100 for the validation set, and 300 for the test set. For WSC, we randomly sample 50 examples as the training set, 50 for the validation set, and 150 for the test set.
D.2 Implementation Details
For the task model, we use both the base model (i.e., Llama-2-7b) and the instruction-tuned models (i.e., Llama-2-7b-chat, GPT-3.5-turbo, and GPT-4). The temperature is 0. For the prompt optimizer, we use GPT-3.5-turbo and GPT-4. The temperature is 1.0. Unless otherwise specified, we use Llama-2-7b-chat as the task model and GPT-3.5-turbo as the prompt optimizer throughout experiments. In the meta-prompt, we include 3 task examples and 7 historical task prompts. For the initial task prompt, we use the original ones from Hendrycks et al. (2021) for MMLU, “Let’s think step by step.” from Kojima et al. (2022) for GSM8K and BBH, and “Let’s solve the problem.” from Yang et al. (2023) for WSC and WebNLG. At each step, the optimizer generates 8 candidates, and the best-performing one is selected. The optimization lasts for at most 3 epochs with a batch size of 8. We repeat all the experiments three times and report the average of the results. The meta-prompts are detailed in Appendix E.2.
Appendix E Meta-Prompt
E.1 Analogical Analysis
Here are the meta-prompts we used in Section 2.
$\bullet$ Prompt+performance.
Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score}
$\bullet$ Prompt+performance+reflection
Your task is to point out the problems with the current prompt based on the wrong examples. The current prompt is: {current prompt} But this prompt gets the following examples wrong. You should analyze the differences between wrong predictions and ground truth answers, and carefully consider why this prompt led to incorrect predictions. Below are the task examples with Queston, Wrong prediction, and Ground truth answer. {error demonstrations} Give a reason why the prompt could have gotten these examples wrong. Wrap the reason with <START> and <END>.
$\bullet$ Summarization-based trajectory
Your task is to integrate the problems in the previous prompt and the current prompt. Below are the problems that arose from the previous prompts. {previous problems} Below are the problems of the current prompt. {current problem} You should integrate the problems of the previous prompt and the current prompt. Wrap the integrated problems with <START> and <END>.
$\bullet$ Retrieval-based trajectory
Below are the previous prompts with their scores. The score ranges from 0 to 100, and higher scores indicate better quality. Prompt: {prompt1} Score: {score1} Prompt: {prompt2} Score: {score2} Prompt: {prompt3} Score: {score3}
$\bullet$ Editing-based refinement
Your task is to modify the current prompt to replace <Prompt>. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace <Prompt> in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Modify the current prompt and get a new improved prompt to replace <Prompt> {prompt position description} in the task examples. Wrap the modified prompt with <START> and <END>.
$\bullet$ Generation-based refinement
Your task is to write a prompt to replace <Prompt>. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace <Prompt> in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Write a new improved prompt to replace <Prompt> {prompt position description} in the task examples. Wrap the new prompt with <START> and <END>.
E.2 Experiment
Here are the meta-prompts we used in Section 4.
$\bullet$ SGDM
The meta-prompt for gradient
Your task is to point out the problems with the current prompt based on the wrong examples. The current prompt is: {current prompt} But this prompt gets the following examples wrong. You should analyze the differences between wrong predictions and ground truth answers, and carefully consider why this prompt led to incorrect predictions. Below are the task examples with Queston, Wrong prediction, and Ground truth answer. {error demonstrations} Give a reason why the prompt could have gotten these examples wrong. Wrap the reason with <START> and <END>.
The meta-prompt for momentum
Your task is to integrate the problems in the previous prompt and the current prompt. Below are the problems that arose from the previous prompts. {previous problems} Below are the problems of the current prompt. {current problem} You should integrate the problems of the previous prompt and the current prompt. Wrap the integrated problems with <START> and <END>.
The meta-prompt for update
Your task is to modify the current prompt to replace <Prompt>. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} Below are the problems with this prompt. {problems} The following exemplars show how to apply the prompt: you replace <Prompt> in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Modify the current prompt and get a new improved prompt to replace <Prompt> {prompt position description} in the task examples. Wrap the modified prompt with <START> and <END>.
$\bullet$ APE
The meta-prompt for update
Your task is to write a prompt to replace <Prompt>. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace <Prompt> in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Write a new improved prompt to replace <Prompt> {prompt position description} in the task examples. Wrap the new prompt with <START> and <END>.
$\bullet$ APO
The meta-prompt for gradient
Your task is to point out the problems with the current prompt based on the wrong examples. The current prompt is: {current prompt} But this prompt gets the following examples wrong. You should analyze the differences between wrong predictions and ground truth answers, and carefully consider why this prompt led to incorrect predictions. Below are the task examples with Queston, Wrong prediction, and Ground truth answer. {error demonstrations} Give a reason why the prompt could have gotten these examples wrong. Wrap the reason with <START> and <END>.
The meta-prompt for update
Your task is to modify the current prompt to replace <Prompt>. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} Below are the problems with this prompt. {problems} The following exemplars show how to apply the prompt: you replace <Prompt> in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Modify the current prompt and get a new improved prompt to replace <Prompt> {prompt position description} in the task examples. Wrap the modified prompt with <START> and <END>.
$\bullet$ OPRO
The meta-prompt for update
Your task is to write a prompt to replace <Prompt>. Below are the previous prompts with their scores. The score ranges from 0 to 100, and higher scores indicate better quality. Prompt: {prompt1} Score: {score1} Prompt: {prompt2} Score: {score2} Prompt: {prompt3} Score: {score3} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace <Prompt> in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Carefully analyze the previous prompts and their scores, and write a new improved prompt to replace <Prompt> {prompt position description} in the task examples. Wrap the new prompt with <START> and <END>.
$\bullet$ PE2
The meta-prompt for gradient
Your task is to point out the problems with the current prompt based on the wrong examples. The current prompt is: {current prompt} But this prompt gets the following examples wrong. You should analyze the differences between wrong predictions and ground truth answers, and carefully consider why this prompt led to incorrect predictions. Below are the task examples with Queston, Wrong prediction, and Ground truth answer. {error demonstrations} Give a reason why the prompt could have gotten these examples wrong. Wrap the reason with <START> and <END>.
The meta-prompt for update
Your task is to write a prompt to replace <Prompt>. Below are the previous prompts with their scores. The score ranges from 0 to 100, and higher scores indicate better quality. Prompt: {prompt1} Score: {score1} Prompt: {prompt2} Score: {score2} Prompt: {prompt3} Score: {score3} The current prompt is: {current prompt} Below are the problems with this prompt. {problems} The following exemplars show how to apply the prompt: you replace <Prompt> in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Carefully analyze the previous prompts and their scores, and write a new improved prompt to replace <Prompt> {prompt position description} in the task examples. You are allowed to change up to {modified word number} words in the current prompt. Wrap the new prompt with <START> and <END>.
$\bullet$ GPO
The meta-prompt for update
Your task is to write a prompt to replace <Prompt>. Below are the previous prompts with their scores. The score ranges from 0 to 100, and higher scores indicate better quality. Prompt: {prompt1} Score: {score1} Prompt: {prompt2} Score: {score2} Prompt: {prompt3} Score: {score3} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace <Prompt> in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Carefully analyze the previous prompts and their scores, and write a new improved prompt to replace <Prompt> {prompt position description} in the task examples. You are allowed to change up to {modified word number} words in the current prompt. Wrap the new prompt with <START> and <END>.
Appendix F Prompts Optimized by Different Methods
Here, we present the prompts optimized by all the methods on the lite BBH benchmark. Prompts on the other tasks can be seen at https://github.com/RUCAIBox/GPO.

### v3

@@ B.3 The Impact of Initial Prompts
In our main experiment, we take “Let’s think step by step.” as the initial prompt. In this part, we aim to explore the impact of initial task prompts. Specifically, we consider prompts from four categories (instructive, misleading, irrelevant, and empty) following Kojima et al. (2022). Table 9 presents the results on the lite BBH benchmark. In general, GPO can boost performance across various initial prompts. This observation underscores the versatility of GPO as a prompt optimizer. Furthermore, the efficacy of optimization is more pronounced with relevant initial prompts (i.e., instructive and misleading). It means that prompt initialization is still important, especially in conveying task-specific information.
B.4 Hyper-Parameter Analysis
@@
B.5 The Execution Time of LLM-based Prompt Optimizers
Figure 6: Execution time of LLM-based prompt optimizers
To further assess the computation costs of LLM-based prompt optimizers, we conduct five repetition experiments and record their average execution time. Figure 6 shows the results on the lite BBH benchmark. We can observe that the execution time is almost consistent with the trend of token consumption. The average execution time of GPO is much lower than SGDM, APO, and PE2, and comparable to OPRO.
C The Interpretability of Optimized Prompts
@@
D Prompts Optimized by Different Methods
In this part, we present the prompts optimized by all the methods on the lite BBH benchmark in Table 12, Table 13, Table 14 and Table 15.
E Detailed Experiment Results
MMLU Tasks Empty CoT SGDM APE APO OPRO PE2 GPO
abstract algebra 27.12 22.03 27.33 32.20 27.12 25.42 30.51 32.20
anatomy 28.33 26.67 23.33 30.00 28.33 31.67 30.00 36.67
astronomy 35.00 36.67 38.33 33.33 35.00 43.33 41.67 41.67
business ethics 37.29 35.59 31.67 27.12 37.29 35.59 45.76 37.29
clinical knowledge 36.67 41.67 43.33 38.33 40.00 38.33 36.67 38.33
college biology 36.67 38.33 45.00 43.33 43.33 45.00 45.00 48.33
college chemistry 37.29 33.90 33.90 35.59 35.59 40.68 37.29 38.98
college computer science 37.29 38.98 33.33 40.68 35.59 38.98 38.98 37.29
college mathematics 30.51 27.12 38.98 30.51 38.98 32.20 32.20 33.90
college medicine 30.00 40.00 31.45 33.33 31.67 28.33 26.67 31.67
college physics 23.33 31.67 36.67 36.67 35.00 33.90 38.98 31.67
computer security 44.07 37.29 44.07 38.98 42.37 38.98 33.90 35.59
conceptual physics 28.33 38.33 35.00 41.67 35.00 36.67 35.00 33.33
econometrics 25.00 28.33 23.33 28.33 26.67 31.67 23.33 26.67
electrical engineering 35.00 31.67 35.00 35.00 31.67 40.00 30.00 40.00
elementary mathematics 33.33 33.33 36.67 33.33 33.33 36.67 33.33 36.67
formal logic 40.00 46.67 35.00 36.67 38.33 45.00 43.33 36.67
global facts 40.68 32.20 35.33 35.59 38.98 35.59 23.73 35.59
high school biology 28.33 43.33 30.00 33.33 26.67 31.67 33.33 31.67
high school chemistry 28.33 26.67 28.67 31.67 31.67 23.33 30.00 35.00
high school computer science 20.34 28.81 23.73 20.34 30.51 20.34 23.73 22.03
high school european history 41.67 48.33 40.00 46.67 41.67 45.00 48.33 48.33
high school geography 48.33 50.00 40.00 45.00 45.00 50.00 51.67 51.67
high school government and politics 45.00 45.00 35.00 43.33 45.00 41.67 50.00 50.00
high school macroeconomics 45.00 30.00 46.67 36.67 46.67 43.33 46.67 48.33
high school mathematics 18.33 21.67 18.33 25.00 18.33 25.00 20.00 23.33
high school microeconomics 31.67 33.33 35.00 35.00 38.33 31.67 35.00 41.67
high school physics 31.67 35.00 30.00 28.33 28.33 35.00 33.33 31.67
high school psychology 48.33 53.33 50.00 55.00 58.33 53.33 55.00 53.33
high school statistics 36.67 35.00 35.00 30.00 36.67 40.00 35.00 36.67
high school us history 41.67 38.33 38.33 38.33 40.00 41.67 40.00 38.33
high school world history 43.33 46.67 41.67 41.67 43.33 45.00 40.00 43.33
human aging 28.33 26.67 38.33 38.33 35.00 30.00 41.67 40.00
human sexuality 30.00 33.33 31.67 38.33 40.00 38.33 45.00 40.00
international law 38.33 40.00 40.00 41.33 38.33 33.33 40.00 38.33
jurisprudence 35.00 40.00 38.67 38.33 33.33 23.33 43.33 38.33
logical fallacies 40.00 30.00 38.33 40.00 40.00 40.00 30.00 35.00
machine learning 30.00 31.67 31.67 36.67 36.67 41.67 40.00 38.33
management 40.00 46.67 40.00 41.67 48.33 41.67 46.67 43.33
marketing 43.33 46.67 38.33 46.67 45.00 40.67 48.67 56.67
medical genetics 40.68 47.46 47.55 45.76 40.68 38.98 42.37 35.59
miscellaneous 43.33 40.00 36.67 36.67 33.33 40.00 36.67 40.00
moral disputes 26.67 25.00 38.33 28.33 21.67 28.33 30.00 30.00
moral scenarios 30.00 11.67 30.00 31.67 45.00 33.33 26.67 31.67
nutrition 45.00 36.67 40.00 48.33 45.00 43.33 48.33 41.67
philosophy 28.33 33.33 45.00 40.00 40.00 50.00 36.67 45.00
prehistory 25.00 28.33 23.33 31.67 26.67 28.33 20.00 28.33
professional accounting 26.67 23.33 38.33 25.00 28.33 45.00 26.67 30.00
professional law 33.33 33.33 33.33 35.00 31.67 35.00 31.67 33.33
professional medicine 45.00 38.33 43.33 43.33 45.00 38.33 40.00 43.33
professional psychology 38.33 35.00 48.33 33.33 35.00 40.00 35.00 43.33
public relations 35.00 36.67 38.33 41.67 41.67 36.78 38.33 40.00
security studies 48.33 46.67 46.67 45.00 56.67 56.67 48.33 50.00
sociology 46.67 45.00 51.67 46.67 48.33 50.00 51.67 55.00
us foreign policy 59.32 55.93 55.93 55.93 57.63 49.15 57.63 59.32
virology 26.67 25.00 30.00 33.33 28.33 26.67 33.33 28.33
world religions 51.67 60.00 56.67 53.33 43.33 55.00 63.33 58.33
Average 35.96 36.36 37.20 37.50 37.71 38.05 38.07 39.14
Table 16: Detailed Results (Accuracy%) on MMLU.
BBH Tasks Empty CoT SGDM APE APO OPRO PE2 GPO
Boolean Expressions 38.89 43.06 48.61 58.33 43.75 42.36 45.14 51.39
Causal Judgement 28.04 23.36 46.73 44.86 42.99 49.53 50.47 47.66
Date Understanding 29.86 36.11 34.03 22.22 35.42 34.72 34.72 33.33
Disambiguation QA 53.47 37.50 46.53 33.33 40.97 29.17 34.03 47.92
Dyck Languages 1.39 2.78 5.56 7.64 2.78 4.86 2.08 4.17
Formal Fallacies 59.03 54.17 55.56 54.86 54.86 48.61 45.83 55.56
Geometric Shapes 6.94 14.58 24.31 22.92 29.17 22.22 27.08 28.47
Hyperbaton 49.31 63.89 63.19 61.81 61.55 75.00 65.97 62.50
Logical Deduction (five objects) 20.83 22.92 25.69 22.92 24.31 23.91 25.00 23.61
Logical Deduction (seven objects) 19.44 17.36 13.89 12.50 13.19 18.75 16.67 17.36
Logical Deduction (three objects) 45.14 36.11 37.50 42.36 34.72 36.81 36.81 41.67
Movie Recommendation 27.08 27.78 48.61 50.69 48.61 52.08 54.17 50.69
Multi-Step Arithmetic 0.00 2.08 0.69 2.78 2.08 0.00 3.47 2.78
Navigate 44.44 43.75 59.03 51.39 49.57 49.84 47.22 51.93
Object Counting 32.64 36.81 36.81 43.75 50.69 44.44 45.83 50.86
Penguins in a Table 31.71 19.51 29.27 26.83 20.73 24.39 20.73 26.83
Reasoning about Colored Objects 29.17 22.92 22.92 21.53 21.53 25.00 21.53 20.83
Ruin Names 27.78 30.56 23.61 30.56 29.17 27.78 29.86 31.94
Salient Translation Error Detection 15.97 13.19 7.64 15.28 18.75 19.44 23.61 18.75
Snarks 46.08 59.80 62.75 64.71 52.94 56.86 59.80 63.73
Sports Understanding 54.86 59.72 59.72 52.08 63.19 59.03 63.19 62.50
Temporal Sequences 20.83 22.92 27.78 19.44 29.17 22.22 25.00 23.61
Tracking Shuffled Objects (five objs) 18.06 10.42 11.81 12.84 10.42 16.67 12.50 17.36
Tracking Shuffled Objects (seven objs) 14.58 9.72 13.19 17.36 10.42 17.36 13.19 16.67
Tracking Shuffled Objects (three objs) 36.11 32.64 30.56 29.86 34.72 29.86 33.33 34.03
Web of Lies 47.22 47.22 45.14 44.44 52.08 45.83 44.44 49.01
Word Sorting 25.00 16.67 18.08 22.22 12.50 22.19 20.83 21.53
All Tasks (avg) 30.51 29.91 33.30 32.94 32.97 33.29 33.43 35.43
Table 17: Detailed Results (Accuracy (%)) on the BBH benchmark.
In this section, we provide the detailed results of the experiment on MMLU in Table 16 and BBH in Table 17.
F Meta-Prompt
F.1 Analogical Analysis
Here are the meta-prompts we used in the Analogical Analysis Section of the main body.
Prompt+performance. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score}
Prompt+performance+reflection Your task is to point out the problems with the current prompt based on the wrong examples. The current prompt is: {current prompt} But this prompt gets the following examples wrong. You should analyze the differences between wrong predictions and ground truth answers, and carefully consider why this prompt led to incorrect predictions. Below are the task examples with Queston, Wrong prediction, and Ground truth answer. {error demonstrations} Give a reason why the prompt could have gotten these examples wrong. Wrap the reason with START and END.
Summarization-based trajectory Your task is to integrate the problems in the previous prompt and the current prompt. Below are the problems that arose from the previous prompts. {previous problems} Below are the problems of the current prompt. {current problem} You should integrate the problems of the previous prompt and the current prompt. Wrap the integrated problems with START and END.
Retrieval-based trajectory Below are the previous prompts with their scores. The score ranges from 0 to 100, and higher scores indicate better quality. Prompt: {prompt1} Score: {score1} Prompt: {prompt2} Score: {score2} Prompt: {prompt3} Score: {score3}
Editing-based refinement Your task is to modify the current prompt to replace Prompt Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace Prompt in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Modify the current prompt and get a new improved prompt to replace Prompt {prompt position description} in the task examples. Wrap the modified prompt with START and END.
Generation-based refinement Your task is to write a prompt to replace Prompt. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace Prompt in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Write a new improved prompt to replace Prompt {prompt position description} in the task examples. Wrap the new prompt with START and END.
F.2 Experiment
Here are the meta-prompts we used in the Experiments Section of the main body.
APE The meta-prompt for update: Your task is to write a prompt to replace Prompt. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace Prompt in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Write a new improved prompt to replace Prompt {prompt position description} in the task examples. Wrap the new prompt with START and END.
APO The meta-prompt for gradient: Your task is to point out the problems with the current prompt based on the wrong examples. The current prompt is: {current prompt} But this prompt gets the following examples wrong. You should analyze the differences between wrong predictions and ground truth answers, and carefully consider why this prompt led to incorrect predictions. Below are the task examples with Queston, Wrong prediction, and Ground truth answer. {error demonstrations} Give a reason why the prompt could have gotten these examples wrong. Wrap the reason with START and END. The meta-prompt for update: Your task is to modify the current prompt to replace Prompt. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} Below are the problems with this prompt. {problems} The following exemplars show how to apply the prompt: you replace Prompt in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Modify the current prompt and get a new improved prompt to replace Prompt {prompt position description} in the task examples. Wrap the modified prompt with START and END.
OPRO The meta-prompt for update: Your task is to write a prompt to replace Prompt. Below are the previous prompts with their scores. The score ranges from 0 to 100, and higher scores indicate better quality. Prompt: {prompt1} Score: {score1} Prompt: {prompt2} Score: {score2} Prompt: {prompt3} Score: {score3} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace Prompt in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Carefully analyze the previous prompts and their scores, and write a new improved prompt to replace Prompt {prompt position description} in the task examples. Wrap the new prompt with START and END.
PE2 The meta-prompt for gradient: Your task is to point out the problems with the current prompt based on the wrong examples. The current prompt is: {current prompt} But this prompt gets the following examples wrong. You should analyze the differences between wrong predictions and ground truth answers, and carefully consider why this prompt led to incorrect predictions. Below are the task examples with Queston, Wrong prediction, and Ground truth answer. {error demonstrations} Give a reason why the prompt could have gotten these examples wrong. Wrap the reason with START and END. The meta-prompt for update: Your task is to write a prompt to replace Prompt. Below are the previous prompts with their scores. The score ranges from 0 to 100, and higher scores indicate better quality. Prompt: {prompt1} Score: {score1} Prompt: {prompt2} Score: {score2} Prompt: {prompt3} Score: {score3} The current prompt is: {current prompt} Below are the problems with this prompt. {problems} The following exemplars show how to apply the prompt: you replace Prompt in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Carefully analyze the previous prompts and their scores, and write a new improved prompt to replace Prompt {prompt position description} in the task examples. You are allowed to change up to {modified word number} words in the current prompt. Wrap the new prompt with START and END.
SGDM The meta-prompt for gradient: Your task is to point out the problems with the current prompt based on the wrong examples. The current prompt is: {current prompt} But this prompt gets the following examples wrong. You should analyze the differences between wrong predictions and ground truth answers, and carefully consider why this prompt led to incorrect predictions. Below are the task examples with Queston, Wrong prediction, and Ground truth answer. {error demonstrations} Give a reason why the prompt could have gotten these examples wrong. Wrap the reason with START and END. The meta-prompt for momentum: Your task is to integrate the problems in the previous prompt and the current prompt. Below are the problems that arose from the previous prompts. {previous problems} Below are the problems of the current prompt. {current problem} You should integrate the problems of the previous prompt and the current prompt. Wrap the integrated problems with START and END. The meta-prompt for update: Your task is to modify the current prompt to replace Prompt. Below is the current prompt with its score. The score ranges from 0 to 100, and higher score indicates better quality. Prompt: {current prompt} Score: {current prompt score} The current prompt is: {current prompt} Below are the problems with this prompt. {problems} The following exemplars show how to apply the prompt: you replace Prompt in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Modify the current prompt and get a new improved prompt to replace Prompt {prompt position description} in the task examples. Wrap the modified prompt with START and END.
GPO The meta-prompt for update: Your task is to write a prompt to replace Prompt. Below are the previous prompts with their scores. The score ranges from 0 to 100, and higher scores indicate better quality. Prompt: {prompt1} Score: {score1} Prompt: {prompt2} Score: {score2} Prompt: {prompt3} Score: {score3} The current prompt is: {current prompt} The following exemplars show how to apply the prompt: you replace Prompt in each input with your new prompt, then read the input and give an output. We say your output is wrong if it is different from the given output, and we say your output is correct if they are the same. {task examples} Carefully analyze the previous prompts and their scores, and write a new improved prompt to replace Prompt {prompt position description} in the task examples. You are allowed to change up to {modified word number} words in the current prompt. Wrap the new prompt with START and END.